[The following information applies to the questions displayed below.] Fitness Fanatics is a regional chain of health clubs. The managers of the clubs, who have authority to make investments as needed, are evaluated based largely on return on investment (ROI). The company's Springfield Club reported the following results for the past year: Sales $ 1,400,000 Net operating income $ 70,000 Average operating assets $ 350,000 The following questions are to be considered independently. 3. Assume that the manager of the club is able to reduce expenses by $14,000 without any change in sales or average operating assets. What would be the club’s return on investment (ROI)?
Answer #1

New net operating income = 70000+14000 = 84000

Return on investment = Operating income/Average operating assets = 84000/350000 = 24%